OpenStudy (anonymous):

x+4y=8 how do i graph this?? how do i get the points?

OpenStudy (anonymous):

it will be easier to graph if you write the same equation in different form: 4y=8-x y=2- x/2 it's a line - you'll need to identify only two points to graph, let's say x=1 (y=2) and... your choice. Finish it

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

i came up with (0,2) (8,0)

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

(0,2) is correct, let's try your second choice: if x=8, y=2- 8/2 = 2-4 = -2 your answer is wrong. it should be (8,-2) for second point

OpenStudy (anonymous):

y=-1/4x + 2 so if x=4 then y=1 x=0 then y = 2 x = 8 then y = 0

OpenStudy (anonymous):

Oups, you are right, I made a mistake in original eq... it should be: y= 2- x/4 so, (8,0) is CORRECT

OpenStudy (anonymous):

oh ok... thank u so much!

OpenStudy (anonymous):

By the way. The graph is a straihgt line going down (the minus of the x) and cutting the y axis in 2.

OpenStudy (anonymous):

how is it a straight line??